OpenStudy (loganator):

How many possible outcomes are there in rolling an 8 sided die twice

OpenStudy (anonymous):

Let A be the event of rolling a 4. P(A) = 1/6 P(A)P(A)=(1/6)(1/6)=1/36 Therefore, the probability of rolling a 4 twice with two rolls of a number cube is 1/36.

OpenStudy (anonymous):

36 basically @taylor12344 is right

OpenStudy (anonymous):

yay lol i feel so accomplishes

OpenStudy (loganator):

thanks

OpenStudy (anonymous):

@LogaNator pm me

OpenStudy (loganator):

please medal you? am I right?

OpenStudy (anonymous):

no plz message me

OpenStudy (loganator):

ok